Find the derivative of the function:


x^5 e^3x + x + 1 / x


a. x^5 e^3x + 5x^4 e^3x + 1 / x^2

b.3x^5 e^3x + 5x^4 e^3x − 1 / x^2

c. x^5 e^3x + 5x^4 e^3x − 1 / x^2

d. 3x^5 e^3x + 5x^4 e^3x − 2 / x






1
Expert's answer
2022-04-22T13:53:06-0400

"\dfrac{d}{dx}(x^5e^{3x}+x+\dfrac{1}{x})=5x^4e^{3x}+3x^5e^{3x}+1-\dfrac{1}{x^2}"